" Using the Section 179 deduction increases the deduction in the initial year, but leaves smaller deductions available in future years. Section 179 does not change the total deduction of the asset over multiple years. It changes the timing of the deduction, putting a larger deduction in the initial year, leaving smaller deductions for later years (Section 179)."

Required

why is it that a business would rather have the larger 179 deduction in the initial year?
Answer #1

Section 179 deduction provides a special deduction of upto $1million in the first year of its purchase of a qualified business property which includes Office Equipment , computers , Machinery and Vehicle weighing more than 6000 pounds.

Section 179 deduction can be useful for a business which purchases new assets on regular basis and have huge taxable incomes.

Section 179 deduction can help them to reduce there tax burden and help them in the cash flows for purchase of new asset for example a hotel or a restaurant. As the hotel keeps on purchasing new assets, it's asset side would keep on increasing without giving any extra benefit in tax cut.

Example- Year 1 Asset purchased $ 1 lakhs and income before section 179 deduction is $ 3 lakhs and Year 2 Asset purchased $50,000 and income of $1.5 Lakhs. Life of Both assets 10 Years and tax rate 30%

With normal depreciation-

Year 1- 3,00,000 - (1,00,000/10) = 2,90,000*0.30 Tax =$ 87,000

Year 2-. 1,50,000 - (1,00,000/10) - (50,000/10) = 1,35,000*0.30 Tax = $40,500

With Section 179 deduction-

Year 1 $3,00,000 - $1,00,000= 2,00,000*0.30Tax = $60,000

Year 2 $1,50,000 - $50,000 = 1,00,000*0.30 Tax = $ 30,000

As we can say from above example in Year 1 $27,000 is left in the cash flow which can find the Purchase of asset in year 2